OpenStudy (anonymous):

The perimeter of an isosceles triangle is 112 inches. The area of the triangle is 588 square inches. a) Find x. Answer in units of inches b) Find y. Answer in units of inches

OpenStudy (nikato):

wait, what is x and y?

OpenStudy (anonymous):

it's number 9 on the pdf file

OpenStudy (nikato):

perimeter: 2y+x=112

OpenStudy (nikato):

area of trinalge= base x heightx 1/2 so 28(1/2)(x)=588 and you can simplify the equation so 14x=588 divide both sides by 14 x=42 substitute this value of x into any one of the orignal 2 equations to solve for y

OpenStudy (anonymous):

so x= 42 y=35?

OpenStudy (nikato):

correct

OpenStudy (anonymous):

can you help me with another problem in the pdf file?

OpenStudy (nikato):

sure. which one?

OpenStudy (anonymous):

number 16

OpenStudy (nikato):

similar to the glove one let d= hot dogs b=burgers

OpenStudy (anonymous):

d+b=8

OpenStudy (nikato):

d+b=8 1.6d+5b=19.6

OpenStudy (anonymous):

to find d I will do 1.6(8-b)+5b=19.6?

OpenStudy (nikato):

no. this will get you to find b but you can still keep going to find b and then plug it back to the equation to find d

OpenStudy (anonymous):

d=6 b=2?

OpenStudy (nikato):

correct agian

OpenStudy (anonymous):

I got it can you help me on number 18?

OpenStudy (nikato):

sorry. i dont know

OpenStudy (anonymous):

number 20?

OpenStudy (nikato):

not sure aboout 20

OpenStudy (anonymous):

22?

OpenStudy (nikato):

no, sorry dont really know much about determinant

OpenStudy (anonymous):

okay thank you for what you have done

OpenStudy (nikato):

your welcome